Grade 11Mechanics

Q.A block starts moving on a horizontal plane under the influence of horizontal acceleration 'a'.After 'to' sec,the acceleration remains same in magnitude but reverses in direction.At what time does it come back to initial point?
Ans.(2+_/2)to

jiten
8 Years agoGrade 11

1 Answer

Karan Gupta
8 Years ago
At time ‘to’, v = u + at = 0 + a’to’ means v = ato. 
Since, acceleration is in opposite direction now, the particle will come to a complete halt in another to, while covering a distance of a total of 2*(1/2*a*to^2) = a(to)^2
 
After covering distance of a(To)^2, particle starts its journey in reverse direction with zero initial velocity.
v^2 = 0 + 2*a*s   [s= a(to)^2]
v^2 = 2*a*(a(to)^2)
v = sqrt(2)*a*to = a*t means t = sqrt (2)*to
 
Hence total time = to + to + sqrt(2)*to
                           =to(2+sqrt(2))